# RepoGraph: Enhancing AI Software Engineering with Repository-level Code Graph

### v1

@@ Abstract
Large Language Models (LLMs) excel in code generation yet struggle with modern AI software engineering tasks. Unlike traditional function-level or file-level coding tasks, AI software engineering requires not only basic coding proficiency but also advanced skills in managing and interacting with code repositories. However, existing methods often overlook the need for repository-level code understanding, which is crucial for accurately grasping the broader context and developing effective solutions. On this basis, we present RepoGraph, a plug-in module that manages a repository-level structure for modern AI software engineering solutions. RepoGraph offers the desired guidance and serves as a repository-wide navigation for AI software engineers. We evaluate RepoGraph on the SWE-bench by plugging it into four different methods of two lines of approaches, where RepoGraph substantially boosts the performance of all systems, leading to a new state-of-the-art among open-source frameworks. Our analyses also demonstrate the extensibility and flexibility of RepoGraph by testing on another repo-level coding benchmark, CrossCodeEval. Our code is available at https://github.com/ozyyshr/RepoGraph.
RepoGraph: Enhancing AI Software Engineering with Repository-level Code Graph
Siru Ouyang1thanks:   This work was done when Siru, Zilin, Zhihan and Mengzhao were interns at Tencent AI Lab, Seattle., Wenhao Yu2, Kaixin Ma2, Zilin Xiao3*, Zhihan Zhang4*, Mengzhao Jia4*, Jiawei Han1, Hongming Zhang2, Dong Yu2 1 University of Illinois Urbana-Champaign 2 Tencent AI Seattle Lab 3 Rice University 4 University of Notre Dame  siruo2@illinois.edu
@@ 1 Introduction
A key step in addressing repository-level tasks is to understand the structure of a repository and identify related code. To achieve this, retrieval-augmented generation (RAG) and its variants (Zhang et al., 2023; Phan et al., 2024; Wu et al., 2024) have been leveraged, in a procedural manner, to retrieve relevant code files across the repository first, providing context for LLMs for further edition. However, indexing at file-level can only identify semantically similar but not genuinely related code snippets. Instead of using RAG, recent approaches like Agentless (Xia et al., 2024) construct a skeletal format for each file, and directly prompt LLMs to identify relevant files and code lines. However, this method still treats code repositories as flat documents (Zhang et al., 2024), which suffers from limitations of repository structure such as the intricate inter-dependencies cross files. An alternative approach is to design agent frameworks (Yang et al., 2024; Wang et al., 2024b), which enables LLMs to interact with repositories using actions. While LLM agents can freely determine the next action based on current observations, without the grasp of global repository structures, they tend to focus narrowly on specific files, resulting in local optimums. Addressing these limitations requires going beyond semantic matching and developing techniques that enable a deeper understanding of the codebase structure. This will allow LLMs to leverage fine-grained context across multiple files and function calls, facilitating more informed, repository-wide decision-making for coding tasks.
@@
Motivated by this, we propose RepoGraph, a plug-in module designed to help LLMs-based AI programmers leverage the code structure of an entire repository. RepoGraph is a graph structure and operates at the line level, offering a more fine-grained approach compared to previous file-level browsing methods. Each node in the graph represents a line of code, and edges represent the dependencies of code definitions and references. RepoGraph is constructed via code line parsing and encodes the structured representation of the entire repository. Sub-graph retrieval algorithms are then used to extract ego-graphs centered around specific keywords. These ego-graphs can be smoothly integrated with both procedural and agent frameworks, offering key clues that provide a more comprehensive context for LLMs to solve real-world software engineering problems.
@@
To assess RepoGraph’s effectiveness and versatility as a plug-in module, we integrate it with four existing software engineering frameworks and evaluate its performance using SWE-bench, a recent benchmark for AI software engineering. Experiment results show that RepoGraph boosts the success rate of existing methods for both agent and procedural frameworks by achieving an average relative improvement of $32.8\%$. We also test RepoGraph on CrossCodeEval to verify its transferibility to general coding tasks that require repository-level code understanding. Additionally, we systematically analyze different sub-graph retrieval algorithms and integration methods. Together with error analyses, we hope to shed light on future works targeting modern AI software engineering.
@@ 2.1 LLM-based methods for AI software engineering
LLM agent-based framework equips large language models (LLMs) with a set of predefined tools, allowing agents to iteratively and autonomously perform actions, observe feedback, and plan future steps (Yang et al., 2024; Zhang et al., 2024; Wang et al., 2024b; Cognition, 2024). While the exact set of tools may vary across different agent frameworks, they typically include capabilities such as opening, writing, or creating files, searching for code lines, running tests, and executing shell commands. To solve a problem, agent-based approaches involve multiple actions, with each subsequent turn depending on the actions taken in previous ones and the feedback received from the environment. For example, SWE-agent (Yang et al., 2024) facilitates interactions with the execution environment by designing a special agent-computer interface (ACI). There are various actions, including “search and navigation”, “file viewer and editor”, and “context management”. Another work, AutoCodeRover (Zhang et al., 2024), further offers fine-grained searching methods for LLM agents in better contexts without an execution process. Specifically, it supports class and function-level code search. OpenDevin (Wang et al., 2024b), initiated after Devin (Cognition, 2024), is a community-driven platform that integrates widely used agent systems. The action space design of OpenDevin is highly flexible, requiring LLM agents to generate code on the fly.
@@ 2.2 Repository-level Coding Capability
While the pre-trained code LLMs mentioned earlier incorporate repository-level information such as file dependencies, tasks at the repository level often involve more intricate call relationships within their context. Recent works like RepoCoder (Zhang et al., 2023) and RepoFuse (Liang et al., 2024) have started integrating Retrieval-Augmented Generation (RAG) modules to harness additional information from repositories. Building on this, subsequent research has focused on embedding repository-level context into their methodologies. For instance, DraCo (Cheng et al., 2024) introduces importing relationships between files, while Aider (Gauthier, 2024) employs PageRank (Page, 1999) to identify the most significant contextual elements. RepoUnderstander (Ma et al., 2024) and CodexGraph (Liu et al., 2024) model code files as a knowledge graph. Despite similarities in representation, methods vary in how they retrieve information from these structures and utilize it for downstream tasks. Table 1 summarizes the differences between these methods and RepoGraph. RepoGraph surpasses previous approaches by effectively integrating context at the line, file, and repository levels.
@@ 3 RepoGraph
Figure 2: An in-depth illustration of (a) the construction, (b) the integration with procedural frameworks, and (c) the integration with agent frameworks of RepoGraph. Given a code repository, we first utilize AST to construct $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{G}$ consists of “reference” and “definition” node, $\mathcal{E}$ includes “invoke” and “contain” relations. The constructed RepoGraph are then used in procedural frameworks by adding sub-retrieval results into each step, and agent frameworks by adding graph retrieval as an additional action “search_repograph”.
@@ 3.1 Construction
Step 1: Code line parsing. We first traverse the entire repository using a top-down approach to identify all code files as candidates for next-step parsing. This is accomplished by filtering based on file extensions, retaining only those with relevant code file suffixes (e.g., .py) while excluding non-essential file types (e.g., .git or requirements.txt), which are noisy and irrelevant for coding tasks. For each code file, we utilize tree-sitter  [^1] https://pypi.org/project/tree-sitter-languages/ to parse the code, leveraging its Abstract Syntax Tree (AST) framework. The AST provides a tree-based representation of the abstract syntactic structure of the source code, enabling the identification of key elements such as functions, classes, variables, types, and other definitions. While recognizing these definitions is crucial, tracing their usage and references throughout the code is equally important. Tree-sitter facilitates this by capturing the definitions and tracking where they are utilized or referenced within the codebase. For example, in figure 2, we not only identify definitions like class Model and its inherent methods but also references like self._validate_input_units(). After processing each line of code with a tree-sitter, we selectively retain lines that involve function calls and dependency relations, discarding extraneous information. Our focus is primarily on the functions and classes, as these represent the core structural components of the code. By concentrating on these elements and their interrelationships, RepoGraph optimizes the analysis process by excluding less significant details, such as individual variables, which tend to be redundant and less relevant for further processing.
@@
Step 2: Project-dependent relation filtering. After the previous parsing step, we obtain code lines with calling and dependency relations. However, not all relations are useful for fixing issues. Specifically, many default and built-in function/class calls could distract from the project-related ones. Therefore, we additionally introduce a filtering process that excludes the repository-independent relations. Two types of such relations exist: (i) global relation refers to Python standard and built-in functions and classes. (ii) local relation are introduced by third-party libraries, which are specific to the current code file. For global relations, we maintain a comprehensive list of methods from standard and built-in libraries, excluding any identified relations from this list. The list is empirically constructed by gathering methods of the builtins library and default methods such as such as “list” and “tuple”. For example, in figure 2, line inputs = len(input) is excluded since “len” is a default method. For local relations, we parse import statements in the code to identify third-party methods that are included, and exclude them accordingly.
@@
Step 3: Graph organization. At this stage, we construct RepoGraph using code lines as the fundamental building units. The graph can be represented as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ represents the set of nodes, with each node corresponding to a line of code, and $\mathcal{E}$ represents the set of edges, capturing the relationships between these code lines. Each node in $\mathcal{V}$ contains attributes to represent its meta-information, such as line_number, file_name, directory, etc. Additionally, we classify each code line as either a “definition" (def) or a “reference" (ref) to a particular module. A “def" node corresponds to the line where a function, class, or variable is initially defined, while a “ref" node indicates a code line where this entity is referenced or invoked elsewhere in the code. Similar to soft links to “def” nodes, “ref” nodes also represent other variations of invoking methods. For example, in Figure 2, the class definition “class Model” would be a “def" node, while any subsequent usages of Model would be “ref" nodes. Each “def" node may have multiple “ref" nodes associated with it, as a single function or class can be referenced in various places throughout the code. We define two types of edges: $\mathcal{E}_{invoke}$ and $\mathcal{E}_{contain}$. The triple $(\mathcal{V}1,\mathcal{E}_{contain},\mathcal{V}_{2})$ denotes that $\mathcal{V}_{1}$ (e.g., a function definition) contains another module $\mathcal{V}_{2}$ (e.g., an internal function or class). The edge $\mathcal{E}_{contain}$ typically connects a “def" node to its internal components. In contrast, $\mathcal{E}_{invoke}$ represents an invocation relationship, usually connecting a “def" node to a “ref" node, where the reference node includes a dependency on the definition.
@@ 3.2 Utility
Integration with procedural framework. In a procedural framework, LLMs are usually prompted in “localization” and “edition” stages with the given repository context and issue description. In this case, we use search_repograph() before both stages, leveraging our RepoGraph to assist in making more informed decisions at each step. For example, in Figure 2, we first include the subgraph of “separability_matrix” for localization, and then use the localized result “Model” to search in the edition stage. To implement the strategy, we flatten the context of retrieved ego-graphs and append it as part of the prompt. As a result, the LLM generation is conditioned on both retrieved ego-graphs and the context provided by baseline methods, helping the model preserve nuances.
@@ 4.1 Setup
Baselines. We integrate RepoGraph with representative methods from both aforementioned research lines. (i) For procedural frameworks, we evaluate the widely used traditional method, RAG (Lewis et al., 2020), as well as Agentless (Xia et al., 2024), an open-source state-of-the-art approach in this direction. For RAG, we follow its initial setting and use BM25 for file-level retrieval. After that, we append the context of RepoGraph after the code files as part of the prompt. Agentless first performs a hierarchical localization in terms of “file-class/function-edits” and then conducts repair based on localization. The context of RepoGraph is inserted in every step of Agentless. (ii) For agent frameworks, we consider SWE-agent (Yang et al., 2024) and AutoCodeRover (Zhang et al., 2024). For both frameworks, we add an additional action “search_repograph” for the LLM agent as described in Section 3. All the choices in the two research lines incorporate GPT-4 and GPT-4o based methods to ensure the generalization. Detailed implementations and prompts used can be found in Appendix A.
Evaluation metrics. We evaluate all methods across two key dimensions: Accuracy and Average Cost. (i) For accuracy, we report the resolve rate and patch application rate. The resolve rate represents the percentage of issues successfully resolved across all data points. An issue is considered resolved if the submitted patch passes all test scripts. To assess the patch application rate, we attempt to apply the generated patches to the repository using the patch program; successful applications contribute to this metric. (ii) To evaluate cost efficiency, we report two metrics: average cost and average tokens, which refer to the inference cost and the number of input/output tokens used when querying the LLMs, respectively.
Configurations. We use the same GPT version as in the baselines in the experiments. We used GPT-4o (2024-05-13) and GPT-4-Turbo (gpt-4-1106-preview) from OpenAI for evaluation and analyses in our experiments. All evaluation processes are performed in a containerized Docker environment  [^3] https://github.com/aorwall/SWE-bench-docker, ensuring stability and reproducibility, made possible through contributions from the open-source community. For plug-in with procedural frameworks, it usually takes around $2$-$3$ hours to finish. For agent frameworks, the inference time is larger, up to around $10$ hours.
@@
SWE-agent GPT-4o 18.33 55 87.00 \$2.51 245,008
    +RepoGraph GPT-4o 20.33 $\uparrow$2.00 61 $\uparrow$6 90.33 $\uparrow$3.33 \$2.69 262,512

### v2

@@ Abstract
Large Language Models (LLMs) excel in code generation yet struggle with modern AI software engineering tasks. Unlike traditional function-level or file-level coding tasks, AI software engineering requires not only basic coding proficiency but also advanced skills in managing and interacting with code repositories. However, existing methods often overlook the need for repository-level code understanding, which is crucial for accurately grasping the broader context and developing effective solutions. On this basis, we present RepoGraph, a plug-in module that manages a repository-level structure for modern AI software engineering solutions. RepoGraph offers the desired guidance and serves as a repository-wide navigation for AI software engineers. We evaluate RepoGraph on the SWE-bench by plugging it into four different methods of two lines of approaches, where RepoGraph substantially boosts the performance of all systems, leading to a new state-of-the-art among open-source frameworks. Our analyses also demonstrate the extensibility and flexibility of RepoGraph by testing on another repo-level coding benchmark, CrossCodeEval. Our code is available at https://github.com/ozyyshr/RepoGraph
@@ 1 Introduction
A key step in addressing repository-level tasks is to understand the structure of a repository and identify related code. To achieve this, retrieval-augmented generation (RAG) and its variants (Xiao et al., 2023; Zhang et al., 2023; Phan et al., 2024; Wu et al., 2024) have been leveraged, in a procedural manner, to retrieve relevant code files across the repository first, providing context for LLMs for further edition. However, indexing at file-level can only identify semantically similar but not genuinely related code snippets. Instead of using RAG, recent approaches like Agentless (Xia et al., 2024) construct a skeletal format for each file, and directly prompt LLMs to identify relevant files and code lines. However, this method still treats code repositories as flat documents (Zhang et al., 2024), which suffers from limitations of repository structure such as the intricate inter-dependencies across files. An alternative approach is to design agent frameworks (Yang et al., 2024; Wang et al., 2024), which enables LLMs to interact with repositories using actions. While LLM agents can freely determine the next action based on current observations, without the grasp of global repository structures, they tend to focus narrowly on specific files, resulting in local optimums. Addressing these limitations requires going beyond semantic matching and developing techniques that enable a deeper understanding of the codebase structure. This will allow LLMs to leverage fine-grained context across multiple files and function calls, facilitating more informed, repository-wide decision-making for coding tasks.
@@
Motivated by this, we propose RepoGraph, a plug-in module designed to help LLMs-based AI programmers leverage the code structure of an entire repository. RepoGraph is a graph structure and operates at the line level, offering a more fine-grained approach compared to previous file-level browsing methods. Each node in the graph represents a line of code, and edges represent the dependencies of code definitions and references. RepoGraph is constructed via code line parsing and encodes the structured representation of the entire repository. Sub-graph retrieval algorithms are then used to extract ego-graphs, which represent the relationships of a central node (in our case, specific keywords). These ego-graphs can be smoothly integrated with both procedural and agent frameworks, offering key clues that provide a more comprehensive context for LLMs to solve real-world software engineering problems.
@@
To assess RepoGraph’s effectiveness and versatility as a plug-in module, we integrate it with four existing software engineering frameworks and evaluate its performance using SWE-bench, a recent benchmark for AI software engineering. Experiment results show that RepoGraph boosts the success rate of existing methods for both agent and procedural frameworks by achieving an average relative improvement of $32.8\%$. We also test RepoGraph on CrossCodeEval to verify its transferability to general coding tasks that require repository-level code understanding. Additionally, we systematically analyze different sub-graph retrieval algorithms and integration methods. Together with error analyses, we hope to shed light on future works targeting modern AI software engineering.
@@ 2.1 LLM-based methods for AI software engineering
LLM agent-based framework equips large language models (LLMs) with a set of predefined tools, allowing agents to iteratively and autonomously perform actions, observe feedback, and plan future steps (Yang et al., 2024; Zhang et al., 2024; Wang et al., 2024; Cognition, 2024; Ouyang et al., 2024; Tang et al., 2025). While the exact set of tools may vary across different agent frameworks, they typically include capabilities such as opening, writing, or creating files, searching for code lines, running tests, and executing shell commands. To solve a problem, agent-based approaches involve multiple actions, with each subsequent turn depending on the actions taken in previous ones and the feedback received from the environment. For example, SWE-agent (Yang et al., 2024) facilitates interactions with the execution environment by designing a special agent-computer interface (ACI). There are various actions, including “search and navigation”, “file viewer and editor”, and “context management”. Another work, AutoCodeRover (Zhang et al., 2024), further offers fine-grained searching methods for LLM agents in better contexts without an execution process. Specifically, it supports class and function-level code search. OpenDevin (Wang et al., 2024), initiated after Devin (Cognition, 2024), is a community-driven platform that integrates widely used agent systems and benchmarks. The action space design of OpenDevin is highly flexible, requiring LLM agents to generate code on the fly.
@@ 2.2 Repository-level Coding Capability
Long before the LLM era, the software engineering community has been studying the navigation of code repositories, including using eye-tracking data (Busjahn et al., 2015) and exploring interconnections (Begel et al., 2010). Then pre-trained code LLMs incorporate repository-level information such as file dependencies. But tasks at the repository level often involve more intricate call relationships within their context. Recent works like RepoCoder (Zhang et al., 2023) and RepoFuse (Liang et al., 2024) have started integrating RAG modules to harness additional information from repositories. Building on this, subsequent research has focused on embedding repository-level context into their methodologies. For instance, DraCo (Cheng et al., 2024) introduces importing relationships between files, CodePlan maintains the code changes by LLMs with incremental dependency analysis, while Aider (Gauthier, 2024) employs PageRank (Page, 1999) to identify the most significant contextual elements. RepoUnderstander (Ma et al., 2024) and CodexGraph (Liu et al., 2024) model code files as a knowledge graph. Despite similarities in representation, methods vary in how they retrieve information from these structures and utilize it for downstream tasks. Table 1 summarizes the differences between these methods and RepoGraph. RepoGraph surpasses previous approaches by effectively integrating context at the line, file, and repository levels.
@@ 3 RepoGraph
Figure 2: An in-depth illustration of (a) the construction, (b) the integration with procedural frameworks, and (c) the integration with agent frameworks of RepoGraph. Given a code repository, we first utilize AST to construct $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{G}$ consists of “reference” and “definition” node, $\mathcal{E}$ includes “invoke” and “contain” relations (files and code lines shown in corresponding color). The constructed RepoGraph are then used in procedural frameworks by adding sub-retrieval results into each step, and agent frameworks by adding graph retrieval as an additional action “search_repograph”. A simplified version can be found in Figure 10.
@@ 3.1 Construction
Step 1: Code line parsing. We first traverse the entire repository using a top-down approach to identify all code files as candidates for next-step parsing. This is accomplished by filtering based on file extensions, retaining only those with relevant code file suffixes (e.g., .py) while excluding other file types (e.g., .git or requirements.txt), which might be noisy and irrelevant for coding tasks. For each code file, we utilize tree-sitter  [^1] https://pypi.org/project/tree-sitter-languages/ to parse the code, leveraging its Abstract Syntax Tree (AST) framework. The AST provides a tree-based representation of the abstract syntactic structure of the source code, enabling the identification of key elements such as functions, classes, variables, types, and other definitions. While recognizing these definitions is crucial, tracing their usage and references throughout the code is equally important. Tree-sitter facilitates this by capturing the definitions and tracking where they are utilized or referenced within the codebase. For example, in figure 2, we not only identify definitions like class Model and its inherent methods but also references like self._validate_input_units(). After processing each line of code with a tree-sitter, we selectively retain lines that involve function calls and dependency relations, discarding extraneous information. Our focus is primarily on the functions and classes, as these represent the core structural components of the code. By concentrating on these elements and their interrelationships, RepoGraph optimizes the analysis process by excluding less significant details, such as individual variables, which tend to be redundant and less relevant for further processing.
@@
Step 2: Project-dependent relation filtering. After the previous parsing step, we obtain code lines with calling and dependency relations. However, not all relations are useful for fixing issues. Specifically, many default and built-in function/class calls could distract from the project-related ones. Therefore, we additionally introduce a filtering process that excludes the repository-independent relations. Two types of such relations exist: (i) global relation refers to Python standard and built-in functions and classes. (ii) local relation are introduced by third-party libraries, which are specific to the current code file. For global relations, we maintain a comprehensive list of methods from standard and built-in libraries, excluding any identified relations from this list. The list is empirically constructed by gathering methods of the builtins library and default methods such as “list” and “tuple”. For example, in figure 2, line inputs = len(input) is excluded since “len” is a default method. For local relations, we parse import statements in the code to identify third-party methods that are included, and exclude them accordingly. A detailed illustration of this step can be found in Figure 11.
@@
Step 3: Graph organization. At this stage, we construct RepoGraph using code lines as the fundamental building units. The graph can be represented as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ represents the set of nodes, with each node corresponding to a line of code, and $\mathcal{E}$ represents the set of edges, capturing the relationships between these code lines. Each node in $\mathcal{V}$ contains attributes to represent its meta-information, such as line_number, file_name, directory, etc. Additionally, we classify each code line as either a “definition” (def) or a “reference” (ref) to a particular module. A “def” node corresponds to the line where a function, class, or variable is initially defined, while a “ref” node indicates a code line where this entity is referenced or invoked elsewhere in the code. Similar to soft links to “def” nodes, “ref” nodes also represent other variations of invoking methods. For example, in Figure 2, the class definition “class Model” would be a “def” node, while any subsequent usages of Model would be “ref” nodes. Each “def” node may have multiple “ref” nodes associated with it, as a single function or class can be referenced in various places throughout the code. We define two types of edges: $\mathcal{E}_{invoke}$ and $\mathcal{E}_{contain}$. The triple $(\mathcal{V}1,\mathcal{E}_{contain},\mathcal{V}_{2})$ denotes that $\mathcal{V}_{1}$ (e.g., a function definition) contains another module $\mathcal{V}_{2}$ (e.g., an internal function or class). The edge $\mathcal{E}_{contain}$ typically connects a “def” node to its internal components. In contrast, $\mathcal{E}_{invoke}$ represents an invocation relationship, usually connecting a “def” node to a “ref” node, where the reference node includes a dependency on the definition.
@@ 3.2 Utility
Integration with procedural framework. In a procedural framework, LLMs are usually prompted in “localization” and “editing” stages with the given repository context and issue description. In this case, we use search_repograph() before both stages, leveraging our RepoGraph to assist in making more informed decisions at each step. For example, in Figure 2, we first include the subgraph of “separability_matrix” for localization, and then use the localized result “Model” to search in the edition stage. To implement the strategy, we flatten the context of retrieved ego-graphs and append it as part of the prompt. As a result, the LLM generation is conditioned on both retrieved ego-graphs and the context provided by baseline methods, helping the model preserve nuances.
@@ 4.1 Setup
Baselines. We integrate RepoGraph with representative methods from both aforementioned research lines. (i) For procedural frameworks, we evaluate the widely used traditional method, RAG (Lewis et al., 2020), as well as Agentless (Xia et al., 2024), an open-source state-of-the-art approach in this direction. For RAG, we follow its initial setting and use BM25 for file-level retrieval. After that, we append the context of RepoGraph after the code files as part of the prompt. Agentless first performs a hierarchical localization in terms of “file-class/function-edits” and then conducts repair based on localization. The context of RepoGraph is inserted in every step of Agentless. (ii) For agent frameworks, we consider SWE-agent (Yang et al., 2024) and AutoCodeRover (Zhang et al., 2024). For both frameworks, we add an additional action “search_repograph” for the LLM agent as described in Section 3. All the choices in the two research lines incorporate GPT-4 and GPT-4o-based methods to ensure generalization. Detailed implementations and prompts used can be found in Appendix A.
Evaluation metrics. We evaluate all methods across two key dimensions: Accuracy and Average Cost. (i) For accuracy, we report the resolve rate and patch application rate. The resolve rate represents the percentage of issues successfully resolved across all data points. An issue is considered resolved if the submitted patch passes all test scripts. To assess the patch application rate, we attempt to apply the generated patches to the repository using the patch program, counting only successful applications toward this metric. (ii) To evaluate cost efficiency, we report two metrics: average cost and average tokens, which refer to the inference cost and the number of input/output tokens used when querying the LLMs, respectively.
Configurations. We use the same GPT version as in the baselines in the experiments. We used GPT-4o (2024-05-13) and GPT-4-Turbo (gpt-4-1106-preview) from OpenAI for evaluation and analyses in our experiments. All evaluation processes are performed in a containerized Docker environment  [^3] https://github.com/aorwall/SWE-bench-docker, ensuring stability and reproducibility, made possible through contributions from the open-source community. Plug-ins with procedural frameworks usually take around $2$-$3$ hours to finish. For agent frameworks, the inference time is larger, up to around $10$ hours.
@@
SWE-agent GPT-4o 18.33 55 87.00 \$2.53 498,346
    +RepoGraph GPT-4o 20.33 $\uparrow$2.00 61 $\uparrow$6 90.33 $\uparrow$3.33 \$2.69 518,792